# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ 3 Experimental Setup
3.1 Simulations
We conduct simulations on synthetic data to assess the theoretical guarantees of our approach. We sampled $X\sim\mathcal{N}(0_{d},I_{d})$ and computed the response $Y=\mu(X)+\sigma(X)\cdot\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,1)$. The mean function $\mu(X)$ introduced non-linearity through transformations of input features (involving absolute values and fractional powers with random coefficients); its explicit form, alongside $\sigma(X)$, is detailed in  Appendix D. The sample size was fixed at $n=5000$ and divided into 70-30% training and validation splits. In the univariate case, we generate 100 datasets for $d=1$, and in multivariate, 100 datasets randomly sampled for $d\in\{2,3,20\}$. We then assess the conditional performance as a function of distance from the training data, where test points are randomly generated on the surfaces of spheres with varying radii (see Appendix D for full explanation).
3.2 Datasets
To evaluate our method in real-world scenarios, we apply it to 17 publicly available regression datasets curated by [23, 39], forming one of the largest benchmarks of its kind and spanning diverse domains such as housing, energy, materials science, and healthcare [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. Categorical features are one-hot encoded, with no further preprocessing. To ensure robustness, each dataset is evaluated over 10 random train-validation-test splits (60%-20%-20%). Note that we use our validation set $\mathcal{D}_{\text{val}}$ also as calibration set $\mathcal{D}_{\text{cal}}$ throughout this paper. A summary of the datasets is provided in Appendix E.
@@ 3.3 Baselines
We compare CLEAR against its core components, CQR and PCS ensemble (we also explore other baselines defined in Section C.1, which we only present in the appendices). Notably, the underlying uncertainty estimation models from CQR and PCS are reused within the CLEAR framework. Our CQR follows classical implementation [9] with the difference that in our experiments, we primarily utilize an enhanced variant, termed ALEATORIC, which uses bootstrapping ($b=100$) on $Y_{i}$ and the model selection from PCS. We further improve ALEATORIC in a new baseline called ALEATORIC-R, which models the residuals $Y_{i}-\hat{f}(X_{i})$ instead, using $\hat{f}$ from the corresponding PCS ensemble. The CLEAR method presented uses the uncalibrated aleatoric uncertainty estimate from this ALEATORIC-R and the uncalibrated epistemic uncertainty from PCS. For fairness and cross-comparison, the model type for quantile estimation within the ALEATORIC-R model is determined by the best-chosen model for epistemic uncertainty.
@@
We perform ablation studies by exploring three different variants of base models for PCS (epistemic) and CQR (aleatoric). Our main approach, which we refer to as variant (a), uses a quantile PCS for estimating epistemic uncertainty. We employ a diverse set of models designed to estimate conditional quantiles, namely quantile random forests (QRF) [34], quantile XGBoost (QXGB) [59], and Expectile GAM [60]. Then, the top-performing model ($k=1$) on the validation set is selected and bootstrapped ($b=100$) to generate the epistemic uncertainty estimate $\hat{q}^{\text{epi}}$ and the median $\hat{f}$. CLEAR then combines this bootstrapped $\hat{q}^{\text{epi}}$ with the bootstrapped aleatoric estimate from ALEATORIC-R, ensuring ALEATORIC-R also uses the same selected quantile model. The other two variants are explained in Section C.2, both using the same $b$ and $k$ as above and only modifying the models. Variant (b) restricts quantile PCS as well as CQR baselines to only QXGB to remove any impact on CLEAR’s evaluation due to the choice of the base models. In contrast, variant (c) uses the standard PCS models from [24, 23] to estimate the conditional mean. In all cases, PCS intervals are calibrated using the multiplicative method.
@@
CLEAR parameters ($\lambda,\gamma_{1}$) are optimized via quantile loss on the validation set. $\lambda$ is chosen from a dense grid combining linearly spaced values from 0 to 0.09 and logarithmically spaced values from 0.1 to 100 (totaling over 4000 points), and $\gamma_{1}$ is determined via conformal calibration for the chosen $\lambda$. All intervals are evaluated at 95% nominal coverage. In our benchmarks, we only tackle the model perturbations of PCS and not the data processing part, where we simplify the process to a single dataset (i.e., $N_{1}=1$) as in [23]. However, we show an example in Section 4.3 of how CLEAR can be applied in the $N_{1}>1$ setting from [24, Chapter 13].
@@ 3.4 Metrics
To evaluate the quality of our prediction intervals, we employ interval coverage (PICP), normalized interval width (NIW), average interval score loss and the quantile loss that are common in the interval prediction literature [61, 62, 63, 64]. We also evaluate our work on Normalized Calibrated Interval Width (NCIW), defined as:
@@
Since we only compare methods that were already calibrated on the validation dataset, they all approximately have the correct coverage on the test set, resulting in $c_{\textrm{test-cal}}\approx 1$. Detailed definitions and closed-form expressions for these metrics are given in Section C.3. We primarily discuss NCIW and quantile loss, but provide the results for all the available metrics.
@@ 4.1 Simulations
Figure 2 shows conditional coverage and interval width for the univariate homoskedastic case. While ALEATORIC-R achieve good coverage in high-density regions but under-cover in low-density or extrapolation regions. When aleatoric and epistemic uncertainties are correlated, all methods perform similarly in the data-rich region, but only CLEAR maintains correct conditional coverage throughout, adapting interval width as needed. Additional results for heteroskedastic and multivariate settings in Appendix D show similar findings.
Figure 2: Results for univariate homoskedastic case averaged over 100 simulations: On the left, conditional coverage, and on the right, mean width, for $X=x$. The plot compares CLEAR, PCS, and ALEATORIC-R (CQR with bootstrapping and trained on residuals $Y_{i}-\hat{f}(X_{i})$). The dashed horizontal line represents the target coverage level of 0.9. CLEAR adapts to maintain target coverage across the input space.
@@ 4.2 Real-world data
Figure 3 shows the Normalized Calibrated Interval Width (NCIW) and Quantile Loss for 95% prediction intervals across all datasets for our main configuration, CLEAR (variant a), compared to several baselines. CLEAR consistently demonstrates superior performance, achieving better or comparable interval width and loss metrics while rigorously maintaining nominal coverage.
@@
The inset boxplots in Figure 3 quantitatively illustrate this advantage: compared to CLEAR (a), PCS, ALEATORIC, and ALEATORIC-R show increases in quantile loss of 15.8%, 34.5%, and 9.4%, respectively. Similar relative increases are observed for NCIW, with PCS, ALEATORIC, and ALEATORIC-R exhibiting increases of 17.4%, 28.2%, and 3%. Moreover, CLEAR (a) was in fact the top‐performing method on 15 of the 17 datasets while it stayed the most stable compared to the baselines. These trends hold across our other model variants as well (Appendix E for full results): both CLEAR (b) and CLEAR (c) exhibit very similar relative improvements (Figures 7 and 8 in Sections E.2 and E.3), with variant (a) remaining the strongest and most robust configuration.
@@
Figure 3: Results for real-world data: Quantile loss and NCIW performance of different methods over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. Error bars represent ±$\pm 1$ standard deviation across seeds. The inset boxplot shows the % of the relative increase of the metric over CLEAR aggregated over all datasets, with the printed numbers displaying an average increase. EPISTEMIC is using the PCS-UQ framework [23]. ALEATORIC is an improved CQR (PCS-inspired with bootstrapping according to [24, Chapter 13]). ALEATORIC-R is further improved by our residual approach. All methods use $\hat{f}$ from PCS.
@@
While we observed that setting $\lambda=1$ or $\gamma_{1}=1$ could marginally improve results if one had prior knowledge of uncertainty components—an unlikely scenario in practice—fully optimizing both parameters offers greater robustness. A limited size of the validation dataset can lead to overfitting of the two parameters. Incorporating some prior on these 2 parameters could potentially further improve CLEAR results. Importantly, the absolute value of $\lambda$ is relative to the pre-calibrated scales of the uncertainty estimators and dataset noise; its interpretation is best contextualized by observing its behavior when systematically varying dataset characteristics, such as the number of features or observations.
Although the reliability of the prediction intervals is critical, we noticed that stand-alone baselines, including ALEATORIC and PCS ensembles, can occasionally exhibit instability on certain datasets, potentially resulting in overly wide or miscalibrated prediction intervals. This characteristic is evident in instances such as the energy_efficiency dataset (Table 9), where ALEATORIC shows a markedly larger NCIW compared to CLEAR. In contrast, CLEAR’s residual modeling and dual-parameter calibration mechanism enhance stability by adaptively re-weighting potentially unreliable or mis-scaled uncertainty components, particularly epistemic estimates; CLEAR consistently produces sharper and more reliable intervals. The extensive dataset-dependent variability observed in the optimal $\lambda$ further highlights the necessity of CLEAR’s adaptive $\lambda$ selection over fixed heuristics, enabling performance that is optimally tailored to the specific characteristics of each dataset.
4.3 Case study: Complete PCS framework
@@
We consider a case study on the Ames Housing dataset, detailed in Appendix F, where (1) we demonstrate the full PCS pipeline, capturing uncertainty from both data preprocessing and model evaluation and (2) we vary the number of predictor variables used to model housing prices. Starting from approximately 80 features, we construct a reduced version using only the top two predictors based on feature importance; this setup naturally balances between aleatoric uncertainty (increases due to limited information) and epistemic uncertainty (decreases due to reduced model complexity). Table 1 reveals that while PCS performs better with many predictors and CQR excels with fewer features, CLEAR adapts effectively to both scenarios through its calibrated uncertainty combination. CLEAR estimates $\lambda=0.6$ in the 2-variable case (prioritizing aleatoric uncertainty) versus $\lambda=14.5$ in the full-feature case (emphasizing epistemic uncertainty), with corresponding epistemic-to-aleatoric ratios of 0.03 and 7.72. This adaptive weighting enables CLEAR to maintain sharp and well-calibrated intervals across both scenarios, evidenced by the nearly halved average interval width when using all features; suggesting that adding covariates effectively reduces uncertainty.
@@ 5 Conclusion, Limitations & Future Work
This paper introduces CLEAR, a novel framework for constructing prediction intervals by adaptively balancing epistemic and aleatoric uncertainty. Through a calibration process involving two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, CLEAR offers an improvement over conventional methods that often address these uncertainty types in isolation or rely on their fixed, non-adaptive combination. Our evaluations using CQR for aleatoric uncertainty and PCS for epistemic uncertainty show that interval width and quantile loss improve on both simulated and the 17 real-world datasets. The method consistently achieves improved conditional coverage, notably by adapting interval widths appropriately in extrapolation regions, and yields narrower average interval widths compared to a suite of strong baseline methods, thereby providing more precise and reliable uncertainty estimates.
Limitations remain despite CLEAR’s significant advantages and these warrant further discussion. From a practical standpoint, the computational overhead associated with our method involves training an ensemble for epistemic uncertainty and fitting models for aleatoric uncertainty. We must note that bootstrapping for the aleatoric part is not necessary; however, the reliance on ensemble techniques and bootstrapping for epistemic uncertainty estimation, while effective, can face scalability challenges with large datasets or models. The accuracy of CLEAR is intrinsically linked to the quality of the base estimators for epistemic and aleatoric uncertainty, with its re-weighting being sensitive to the size and representativeness of the calibration set. Epistemic uncertainty, as shown in the case study, must be handled using careful judgment calls, in line with the principles of the PCS framework [31, 24]. The additive combination of scaled uncertainties, while powerful, is also a specific structural choice.
Future work could explore several promising directions. A natural modification of the CLEAR framework is its extension to classification tasks, where disentangled uncertainty estimates may also improve calibration [65]. Additionally, investigating better techniques for choosing the $\lambda$ parameter beyond the performance on a validation set might refine its selection process. Another direction is integrating CLEAR with active learning paradigms, where its uncertainty decomposition could guide more efficient data acquisition [15]. Moreover, exploring alternative, potentially more scalable, approaches for epistemic uncertainty quantification, such as those based on Vine Copulas [66] or other non-ensemble methods, could mitigate current computational limitations and ease the framework’s adoption. Extending CLEAR to time series settings could also be valuable, especially for capturing temporal uncertainty dynamics [67, 68, 69, 70, 71, 72, 73, 74, 75]. Finally, a deeper study into the interpretability of the learned $\lambda$ and $\gamma_{1}$ parameters could provide insights into the dominant sources of uncertainty in various contexts.
@@
We are grateful to Abhineet Agarwal and Michael Xiao for their valuable feedback, helpful discussions, and for sharing early previews and detailed explanations of their code. We also thank Natasa Tagasovska, Jakob Weissteiner, Hanna Wutte, Florian Krach, Markus Chardonnet, Josef Teichmann, Bertrand Charpentier, Janis Postels, Alexander Immer, Erez Buchweitz, João Vitor Romano, Omer Ronen, James-A. Goulet, and Hans Bühler for insightful and stimulating discussions. We are also grateful to Valérie Chavez-Demoulin for her continuous support.

### v2

@@ 3 Experimental Setup
3.1 Data
We conduct experiments using both simulations and real-world data. The synthetic experiments assess the theoretical guarantees of our approach. We sample $X\sim\mathcal{N}(0_{d},I_{d})$ and compute the response $Y=\mu(X)+\sigma(X)\cdot\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,1)$. The mean function $\mu(X)$ introduces non-linearity through transformations of input features (involving absolute values and fractional powers with random coefficients); its explicit form, alongside $\sigma(X)$, is detailed in Appendix C. The sample size is fixed at $n=5000$ and divided into 70-30% training and validation splits. In the univariate case, we generate 100 datasets for $d=1$, and in multivariate, 100 datasets randomly sampled for $d\in\{2,3,20\}$. We then assess the conditional performance as a function of distance from the training data, where test points are randomly generated on the surfaces of spheres with varying radii. For the real-world scenarios, we use 17 regression datasets curated by [30], forming one of the largest benchmarks for UQ (see Section D.1 for details). Categorical features are one-hot encoded, with no further preprocessing. To ensure robustness, each dataset is evaluated over 10 random train-validation-test splits (60%-20%-20%). We conduct experiments in two configurations: standard (using $\mathcal{D}_{\text{val}}$ also as $\mathcal{D}_{\text{cal}}$) and conformalized (splitting the 20% validation into 10% $\mathcal{D}_{\text{val}}$ + 10% $\mathcal{D}_{\text{cal}}$ for stronger theoretical guarantees). Note that while the body of the paper only focuses on the standard experiments, the conformalized results are provided in Appendix G.
@@ 3.2 Baselines
We compare CLEAR against its core components, CQR and PCS ensemble (for details, including PCS implementation and further baselines that are relevant only to the appendices, see Appendix D). Notably, the underlying uncertainty estimation models from CQR and PCS are reused within the CLEAR framework. Our CQR follows the classical implementation [10] with the difference that in our experiments, we primarily utilize an enhanced variant, termed ALEATORIC, which uses bootstrapping ($b=100$) on $Y_{i}$ and the model selection from PCS. We further improve ALEATORIC in a new baseline called ALEATORIC-R, which models the residuals $Y_{i}-\hat{f}(X_{i})$ instead, using $\hat{f}$ from the corresponding PCS ensemble. The CLEAR method presented uses the uncalibrated aleatoric uncertainty estimate from this ALEATORIC-R and the uncalibrated epistemic uncertainty from PCS. For fairness and cross-comparison, the model type for quantile estimation within the ALEATORIC-R model is determined by the best-chosen model for epistemic uncertainty.
@@
We perform ablation studies by exploring three different variants of base models for PCS (epistemic) and CQR (aleatoric). Our main approach, which we refer to as variant (a), uses a quantile PCS for estimating epistemic uncertainty. We employ a diverse set of models designed to estimate conditional quantiles, namely quantile random forests (QRF) [33], quantile XGBoost (QXGB) [38], and Expectile GAM [39]. Then, the top-performing model ($k=1$) on the validation set is selected and bootstrapped ($b=100$) to generate the epistemic uncertainty estimate $\hat{q}^{\text{epi}}$ and the median $\hat{f}$. CLEAR then combines this bootstrapped $\hat{q}^{\text{epi}}$ with the bootstrapped aleatoric estimate from ALEATORIC-R, ensuring ALEATORIC-R also uses the same selected quantile model. The other two variants are explained in Section D.4, both using the same $b$ and $k$ as above and only modifying the models. Variant (b) restricts quantile PCS as well as CQR baselines to only QXGB to remove any impact on CLEAR’s evaluation due to the choice of the base models. In contrast, variant (c) uses the standard PCS models to estimate the conditional mean. In all cases, PCS intervals are calibrated using the multiplicative method.
To further validate the generalizability of CLEAR, we also apply and evaluate it to other uncertainty estimators. This is a separate setup, where for epistemic uncertainty, we employ Deep Ensembles (DE) [22]—an ensemble of neural networks trained with diverse initializations—and for aleatoric uncertainty, we use Simultaneous Quantile Regression (SQR) [13], which directly models multiple conditional quantiles (more detail in Section D.3). These represent state-of-the-art deep learning approaches for uncertainty quantification, complementing our primary PCS and CQR baselines.
@@
In all cases, CLEAR parameters ($\lambda,\gamma_{1}$) are optimized via quantile loss on the validation set. $\lambda$ is chosen from a dense grid combining linearly spaced values from 0 to 0.09 and logarithmically spaced values from 0.1 to 100 (totaling over 4000 points), and $\gamma_{1}$ is determined via conformal calibration for the chosen $\lambda$. All intervals are evaluated at 95% nominal coverage. In our benchmarks, we only address the model perturbations of PCS, not the data processing part, where we simplify the process to a single dataset ($N_{1}=1$), as in [30]. However, we show an example in Section 4.3 of how CLEAR can be applied in the $N_{1}>1$ setting from [40, Chapter 13].
@@ 3.3 Metrics
To evaluate the quality of our prediction intervals, we employ interval coverage (PICP), normalized interval width (NIW), average interval score loss and the quantile loss that are common in the interval prediction literature [41, 42, 43]. We also evaluate our work on Normalized Calibrated Interval Width (NCIW), defined as:
@@
In the standard configuration, all methods are calibrated on the validation set; in the conformalized configuration, calibration uses a separate 10% calibration split. In both cases the methods are already calibrated before evaluation on the test set, resulting in $c_{\textrm{test-cal}}\approx 1$. We primarily discuss NCIW and quantile loss, but provide the results for all the available metrics.
@@ 4.1 Simulations
Figure 2 shows conditional coverage and interval width for the univariate homoskedastic case. While ALEATORIC-R achieve good coverage in high-density regions but under-cover in low-density or extrapolation regions. When aleatoric and epistemic uncertainties are correlated, all methods perform similarly in the data-rich region, but only CLEAR maintains correct conditional coverage throughout, adapting interval width as needed. Additional results for heteroskedastic and multivariate settings in Section C.2.2 show similar findings.
Figure 2: Results for univariate homoskedastic case averaged over 100 simulations: On the left, conditional coverage, and on the right, mean width, for $X=x$. It compares CLEAR, PCS, and ALEATORIC-R (bootstrapped CQR trained on residuals $Y_{i}-\hat{f}(X_{i})$). The dashed horizontal line is the target coverage level of 0.9. CLEAR adapts to maintain target coverage across the input space.
@@ 4.2 Real-world data
Figure 3 shows the Normalized Calibrated Interval Width (NCIW) and Quantile Loss for 95% prediction intervals across all datasets for our main approach, CLEAR (variant a), compared to several baselines. CLEAR consistently demonstrates superior performance, achieving better or comparable interval width and loss metrics while rigorously maintaining nominal coverage. The inset boxplots show that CLEAR (a) compared to PCS, ALEATORIC, and ALEATORIC-R has an improved quantile loss of 15.8%, 34.5%, and 9.4%, respectively. Similar relative increases are observed for NCIW, with PCS, ALEATORIC, and ALEATORIC-R exhibiting increases of 17.4%, 28.2%, and 3%. Moreover, CLEAR (a) was, in fact, the top-performing method on 15 of the 17 datasets, while remaining the most stable compared to the baselines. These trends hold across our other model variants as well (Appendix F for standard and Appendix G for conformalized results): both CLEAR (b) and CLEAR (c) exhibit very similar relative improvements (Figures 7 and 8 in Sections F.2 and F.3), with variant (a) remaining the strongest and most robust configuration.
@@
Figure 3: Results for real-world data: Quantile loss and NCIW performance of different methods over 10 seeds normalized relative to CLEAR (baseline = 1.0) with error bars are $\pm 1$ $\sigma$. Lower values are better. The inset boxplot shows the average (%) relative increase of the metric over CLEAR. EPISTEMIC is PCS-UQ, ALEATORIC is bootstrapped CQR, and ALEATORIC-R uses residuals.
@@
While we observed that setting $\lambda=1$ or $\gamma_{1}=1$ could marginally improve results if one had prior knowledge of uncertainty components—an unlikely scenario in practice—fully optimizing both parameters offers greater robustness. A limited size of the validation dataset can lead to overfitting of the two parameters, and incorporating some prior on them could further improve the results. Importantly, the absolute value of $\lambda$ is relative to the pre-calibrated scales of the uncertainty estimators and dataset noise; its interpretation is best contextualized by observing its behavior when systematically varying dataset characteristics, such as the number of features or observations. While fixed parameters ($\lambda=1$ or $\gamma_{1}=1$) could marginally improve results with prior knowledge, fully optimizing both parameters offers greater robustness. CLEAR’s dual-parameter calibration enhances stability by adaptively re-weighting potentially unreliable uncertainty components, as evidenced in datasets like energy_efficiency, where baselines show markedly larger NCIW. The dataset-dependent variability in optimal $\lambda$ underscores the need for adaptive selection over fixed heuristics. The calibration runtime (grid-search) is also extremely negligible in practice (Section F.5).
Beyond PCS and CQR, CLEAR demonstrates substantial improvements when applied to DE and SQR (Table 1). When using DE for epistemic uncertainty and SQR for aleatoric uncertainty at 95% nominal coverage, CLEAR achieves average width reductions (NCIW) of 28.6% and 13.4%, respectively, with similar improvements in quantile loss (24.0% and 13.7%). The gains persist even after conformal calibration of the baselines, particularly relevant to the aleatoric component (SQR). The results underscore CLEAR’s ability to balance the two uncertainty sources well, regardless of the underlying estimators. This consistency across both PCS and neural approaches validates the generality of CLEAR (full results in Appendix E).
Table 1: Mean (%) improvement of CLEAR over DE & SQR across 17 datasets (higher is better).
Metric DE SQR DE-conformal SQR-conformal
PICP +0.05% -0.66% -0.09% -0.15%
NIW +28.81% +17.38% +29.55% +14.07%
NCIW +28.57% +13.36% +27.90% +13.23%
QuantileLoss +23.98% +13.66% +24.08% +10.12%
4.3 Case study: Accounting For Data Uncertainty
@@
We consider a case study on the Ames Housing dataset, detailed in Appendix H, where we demonstrate the full PCS pipeline and vary the number of predictor variables used to model housing prices. Starting from approximately 80 features, we construct a reduced version using only the top two predictors; this setup naturally balances aleatoric uncertainty (increases due to limited information) and epistemic uncertainty (decreases due to reduced model complexity). Table 2 reveals that while PCS performs better with many predictors and CQR excels with fewer features, CLEAR adapts effectively to both scenarios through its calibrated uncertainty combination. CLEAR estimates $\lambda=0.6$ in the 2-variable case (prioritizing aleatoric uncertainty) versus $\lambda=14.5$ in the full-feature case (emphasizing epistemic uncertainty), with corresponding epistemic-to-aleatoric ratios of 0.03 and 7.72. This adaptive weighting enables CLEAR to maintain sharp intervals across both scenarios, with nearly halved average interval width when using all features.
@@ 5 Conclusion, Limitations & Future Work
This paper introduces CLEAR, a novel framework for constructing prediction intervals by adaptively balancing epistemic and aleatoric uncertainty. Through a calibration process involving two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, CLEAR offers an improvement over classical methods that often address these uncertainty types in isolation or rely on their fixed, non-adaptive combination. Our evaluations using CQR (and SQR) for aleatoric uncertainty and PCS (and DE) for epistemic uncertainty show that interval width and quantile loss improve on both simulated and the 17 real-world datasets. CLEAR consistently achieves improved conditional coverage, notably by adapting interval widths appropriately in extrapolation regions, while yielding narrower interval widths.
Limitations remain despite CLEAR’s significant advantages, and these warrant further discussion. When scaling CLEAR to significantly larger datasets and models, specific hyperparameters (e.g., number of bootstraps) can be adjusted to save computational costs. The accuracy of CLEAR is intrinsically linked to the quality of the base estimators for epistemic and aleatoric uncertainty. All our calibration datasets contained at least 150 data points, but for smaller calibration datasets, overfitting $\gamma_{1}$ and $\lambda$ could be a problem. Epistemic uncertainty, as demonstrated in the case study, must be addressed through careful judgment calls, in line with the principles of the PCS framework. The additive combination of scaled uncertainties, while powerful, is also a specific structural choice.
Future work could explore extension to classification tasks [44], alternative $\lambda$ selection techniques, integration with active learning, and more scalable epistemic UQ approaches [45]. Extending CLEAR to time series settings could also be valuable, particularly for capturing the dynamics of temporal uncertainty. Finally, a deeper study of the interpretability of learned parameters could provide insights into the dominance of the two uncertainty sources.
@@
We are grateful to Abhineet Agarwal and Michael Xiao for their valuable feedback, helpful discussions, and for sharing early previews and detailed explanations of their code. We also thank Natasa Tagasovska, João Vitor Romano, Erez Buchweitz, Omer Ronen, Jakob Weissteiner, Hanna Wutte, Florian Krach, Markus Chardonnet, Josef Teichmann, Bertrand Charpentier, Janis Postels, Alexander Immer, James-A. Goulet, and Hans Bühler for insightful and stimulating discussions. We are also grateful to Valérie Chavez-Demoulin for her continuous support.
@@
Ilia Azizi acknowledges support from the HEC Research Fund grant that enabled and facilitated computations on DCSR clusters at the University of Lausanne.